OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − x^2 = −3x − 10?

OpenStudy (anonymous):

(2, 0) and (−5, 0) (−2, 0) and (5, 0) (−2, 0) and (−5, 0) (2, 0) and (5, 0)

OpenStudy (nathan917):

The x intercept is when the graph of the equation crosses the x-axis, i.e. the Y-coordinate equals zero. So just substitute zero for y and solve for x. 0 = x^2 -3x-10. this is factorable so it is easy and the answers are integers. 0 = (x-5)(x+2) either (x-5) equals 0 or (x+2) equals zero (or both) so if x-5 = 0 x=5 and if x+2 =0 x = -2 Answer: x = -2 and 5. We already know that y=0, so the coordinates for the x-intercepts of the equation above are (-2,0) and (5,0).

OpenStudy (anonymous):

oh ok thnx